eclipsedstar (eclipsedstar):

The average (arithmetic mean) of 4 different integers is 75. If the largest integer is 90, what is the least possible value of the smallest integer? A. 19 B. 29 C. 30 D. 33

OpenStudy (fibonaccichick666):

this is not a nice question. So before I spout into something, what level math is this for?

OpenStudy (kainui):

Set up some equations describing this. That tends to make word problems a little more transparent.

eclipsedstar (eclipsedstar):

I'm in Algebra 2...I'm not sure really, this is a extra credit question my teacher came up with but I don't know how to start at all.

OpenStudy (fibonaccichick666):

ohhhk. Well then

OpenStudy (anonymous):

If the biggest number is 90, how big can the next two largest numbers be?

OpenStudy (fibonaccichick666):

That kicks my number theory approach out of here.

OpenStudy (fibonaccichick666):

I would say, let's set up a few things we know as Kainui recommended

OpenStudy (fibonaccichick666):

so how do we find the mean of 4 numbers?

OpenStudy (kainui):

Well first off let's just give the variables names: \[\Large a,b,c,d\] What do we know about them? We know their average is equal to 75 right? \[\Large \frac{a+b+c+d}{4}=75\] What else do we know? We know they're different integers right? So that means we can just write them as \[\Large a<b<c<d\] What else do we know? We know the largest one is 90, so that must be d. \[\Large d=90\] What can we do from here?

OpenStudy (kainui):

We can now say that we're looking for "a" since that will be the smallest right? So how can we do that? Let's maximize b and c. So b and c equal 88 and 89 since that's the largest they can be. Plug them all in and we can solve for a! Done!

OpenStudy (fibonaccichick666):

Nice! but we also must take into account that it needs to be divisible by 4

OpenStudy (fibonaccichick666):

which we will if we solve for a :)

OpenStudy (fibonaccichick666):

Do you follow @EclipsedStar ?

OpenStudy (kainui):

(I might be late posting this, but I'll post it anyways) I don't think we do, do we? For the more general case we are looking at: \[\Large \frac{a+b+c+d}{4}=t \\ \Large a+b+c+d = 4t \\ \Large a= 4t-b-c-d\] So we don't need a to be divisible by 4.

OpenStudy (fibonaccichick666):

I meant that t needs to be divisible by a, so it can't be even. (so we could get rid of an answer)

eclipsedstar (eclipsedstar):

The answer choices also don't add up. ^

OpenStudy (anonymous):

There's no need for any equations. It's as simple as: What is the sum of the four numbers? What is the biggest we can make the other three numbers? What does that leave for the smallest number?

OpenStudy (fibonaccichick666):

let's see: \[a+88+89+90=4(75)\]solve this @EclipsedStar

OpenStudy (fibonaccichick666):

The answer is there

OpenStudy (kainui):

That's what I've been saying... #_#

OpenStudy (fibonaccichick666):

I know

OpenStudy (fibonaccichick666):

You have my best response for it Kainui

eclipsedstar (eclipsedstar):

33. XD

eclipsedstar (eclipsedstar):

D! :'D

OpenStudy (kainui):

Haha, I just mean I was trying to avoid just sort of giving the answer too much. Oh well

eclipsedstar (eclipsedstar):

Thanks! I get it now. :)

OpenStudy (fibonaccichick666):

True, I should have been less to the point

OpenStudy (fibonaccichick666):

My bad, sorry @Kainui